Responses/last dep 19756724 Your last submission is uned for your score and frs terc a wnth sitis a ene a d benifer exent troand ,on the couh. Andreas force i due (a) Find the net force (in N) in component form 12 (b) Find the magnitude (in N) and direction (in degrees counterclockwise from the east axis) of the net force direction counterclockwise from the east axis tf Andrea and Jennifers housemates, David and Stephanie, disagree with the move and want to prevent its relocation, with what combined force Fos (in N) should they push o that the couch does not move? (Express your answer in vector form) Additional Materials eBook 76 Type here to search
Answer #1

(a)

Andrea's force, FA = 120 j N

Jennifer's force, FJ = 190*sin22 i + 190*cos22 j

FJ = 71.17 i + 176.16 j N

Net force = 71.17 i + (120 + 176.16) j

Net force = 71.17 i + 296.16 j

(b)

Magnitude of net force,

Fnet = sqrt [(71.17)^2 + (296.16)^2]

Fnet = 304.6 N

Direction,

tan\theta = 296.16 / 71.17

\theta = 76.5 deg

(c)

To make the couch not to move, net force on couch = 0

so, FDS = -Fnet

FDS =  (-71.17 i - 296.16 j) N